# Alternating Gradient-Type Algorithm for Bilevel Optimization with Inexact Lower-Level Solutions via Moreau Envelope-based Reformulation

### v2

@@ 1 Introduction
where $\gamma>0$. When the lower-level problem is convex with respect to $y$, this reformulation is equivalent to the original bilevel optimization problem (1) (see [31, Theorem 1]). Directional optimality conditions for $({\rm VP})_{\gamma}$ are discussed in [5]. However, as shown in [31] and [61, Proposition 7], classical constraint qualification (CQ) conditions do not hold at any feasible point of $({\rm VP})_{\gamma}$, motivating us to study the following relaxed approximation problem,
@@ 2.1 Notations
[…]
^$\hat{\partial}h(\bar{z})$ =$=\left\{v\in\mathbb{R}^{n}\;\middle|\;h(z)\geq h(\bar{z})+\langle v,z-\bar{z}\rangle+o(\|z-\bar{z}\|),\;\forall z\in\mathbb{R}^{n}\right\},$
@@ 2.2 Basic Assumptions
Remark 1.
The bilevel problem (1) has also been studied in [42]. However, in addition to Assumption 3, Assumption 3.2 in [42] imposes an additional Lipschitz-like continuity condition on the proximal operator $\operatorname{Prox}_{{g}(x,\cdot)}$ with respect to $x$. This assumption is challenging to verify in practice and is not required in this work.
Assumption 3(1) implies $(\rho_{f_{1}},\rho_{f_{2}})$-weak convexity with $\rho_{f_{1}}=\rho_{f_{2}}=L_{f}$ of $f(x,y)$ [7, Lemma 5.7]. Furthermore, the lower-level objective $\varphi(x,y)$ is $(\rho_{\varphi_{1}},\rho_{\varphi_{2}})$-weakly convex on $X\times Y$, where $\rho_{\varphi_{1}}=\rho_{f_{1}}+\rho_{g_{1}}$ and $\rho_{\varphi_{2}}=\rho_{f_{2}}+\rho_{g_{2}}$.
According to [24, Lemma 4.2], $g(x,y)=h(c(x,y))$ is $L_{h}L_{c}$-weakly convex if $h:\mathbb{R}^{d}\to\mathbb{R}$ is a closed convex $L_{h}$-Lipschitz continuous function and $c:\mathbb{R}^{n+m}\to\mathbb{R}^{d}$ is a smooth function with an $L_{c}$-Lipschitz continuous Jacobian. Below, we provide more specific examples of $g(x,y)$ satisfying Assumption 3.
Example 1.
Consider $g(x,y)=xP(y)$, where $P(y)$ is a convex and $L_{P}$-Lipschitz continuous function on $Y$. Then $g(x,y)$ is $(\rho_{g_{1}},\rho_{g_{2}})$-weakly convex on +×$\mathbb{R}_{+}\times\mathbb{R}^{m}$ for any $\rho_{g_{1}},\rho_{g_{2}}\geq 0$ satisfying $\rho_{g_{1}}\rho_{g_{2}}\geq L_{P}^{2}$.
@@
(1) Let $g(x,y)=x\|y\|_{1}$. Then $g(x,y)=xP(y)$ with $P(y)=\|y\|_{1}$, which is $\sqrt{m}$-Lipschitz on $\mathbb{R}^{m}$. Thus, $g(x,y)$ is $(m,1)$-weakly convex on +×$\mathbb{R}_{+}\times\mathbb{R}^{m}$.
(2) Let $g(x,y)=\sum_{j=1}^{J}x_{j}\|y^{(j)}\|_{2}$, where $y^{(j)}$ denotes the $j$-th group of $y$. Then $g(x,y)=\sum_{j=1}^{J}x_{j}P_{j}(y^{(j)})$ with $P_{j}(y^{(j)})=\|y^{(j)}\|_{2}$, which is $1$-Lipschitz. Thus, $g(x,y)$ is $(1,1)$-weakly convex on +$\mathbb{R}^{J}_{+}\times\mathbb{R}^{m}$.
@@ 2.3 On the Moreau Envelope Function Reformulation
Under Assumption 2, the solution mapping
@@
Proposition 1.
Suppose Assumptions 2 and 3 hold. For $\gamma\in(0,1/(2\rho_{f_{2}}+2\rho_{g_{2}}))$, the function $v_{\gamma}(x,y)$ is $(\rho_{v_{1}},\rho_{v_{2}})$-weakly convex on $X\times\mathbb{R}^{m}$ with $\rho_{v_{1}}\geq\rho_{f_{1}}+\rho_{g_{1}}$ and $\rho_{v_{2}}\geq 1/\gamma$. Additionally, $v_{\gamma}(x,y)$ is differentiable on $X\times\mathbb{R}^{m}$ with gradient
@@
Proposition 2.
Suppose Assumptions 2 and 3 hold. For $\gamma\in(0,1/(\rho_{f_{2}}+\rho_{g_{2}}))$, the function $v_{\gamma}(x,y)$ is $\rho_{v_{1}}$-weakly convex with respect to $x$ on $X$ for any $y\in Y$ with $\rho_{v_{1}}\geq\rho_{f_{1}}+\rho_{g_{1}}$. It is also convex with respect to $y$ on $\mathbb{R}^{m}$ for any $x\in X$. Moreover, $v_{\gamma}(x,y)$ is differentiable with gradient given by (9) on $X\times\mathbb{R}^{m}$, and $\theta_{\gamma}^{*}(x,y)$ and $\nabla v_{\gamma}(x,y)$ are continuous on $X\times\mathbb{R}^{m}$.
@@
We begin by showing that $v_{\gamma}(x,y)$ is continuously differentiable on $X\times\mathbb{R}^{m}$. By Assumption 2, the function $f(x,\theta)+g(x,\theta)+\frac{1}{2\gamma}\|\theta-y\|^{2}$ is level-bounded in $\theta$ locally uniformly for any $(\bar{x},\bar{y})\in X\times\mathbb{R}^{m}$. Specifically, for any $c\in\mathbb{R}$, there exist a compact set $D$ and a neighborhood $Z$ of $(\bar{x},\bar{y})$ such that the level set $\{\theta\in\mathbb{R}^{m}\mid f(x,\theta)+g(x,\theta)+\frac{1}{2\gamma}\|\theta-y\|^{2}\leq c\}$ is contained in $D$ for all $(x,y)\in Z$. Using Corollary 4.3(ii) from [49], it follows that $v_{\gamma}(x,y)$ is locally Lipschitz continuous on $X\times\mathbb{R}^{m}$. Furthermore, Theorem 1.22, Theorem 4.1(ii) and Theorem 4.17 in [49] imply that $v_{\gamma}({x},{y})$ is differentiable on $X\times\mathbb{R}^{m}$ with its gradient given as in (9). The partial weak convexity property of $v_{\gamma}({x},{y})$ can be established using arguments similar to those in the proof of Theorem 2 in [31].
@@
Lemma 1.
Suppose Assumptions 2 and 3 hold. The mapping $\theta_{\gamma}^{*}(x,y)$ is $L_{\theta^{*}}$-Lipschitz continuous with respect to $y$ for any $x\in X$, with $L_{\theta^{*}}:=1+\sqrt{1-s/\gamma}$ where $s=\gamma/(\gamma L_{f_{y}}+1)^{2}$.
@@
$20\in\nabla_{y}f(x,\theta^{*}_{i})+\partial_{y}g(x,\theta^{*}_{i})+(\theta^{*}_{i}-y_{i})/\gamma+\mathcal{N}_{Y}(\theta^{*}_{i}),\quad\mathrm{for}~{}i=1,$ 2 .
@@ 3 Algorithm design
where $p_{k}$ is the penalty parameter at iteration $k$. Due to the nonsmoothness of $g(x,y)$ with respect to $y$ and the availability of its proximal operator, it is natural to adopt a proximal alternating linearization method [14] to construct the algorithm. Given the current iterate $(x^{k},y^{k})$, linearizing the smooth components of $\psi_{p_{k}}(x,y)$ requires evaluating gradients, including $\nabla v_{\gamma}(x,y)$ as provided in Proposition 1. Calculating the exact gradient $\nabla v_{\gamma}(x^{k},y^{k})$ requires solving the proximal lower-level problem in (8) to obtain $\theta_{\gamma}^{*}(x^{k},y^{k})$. However, this solution typically does not have a closed-form expression, and solving it requires an iterative optimization solver, which can be computationally expensive. To address this, we propose an inexact gradient approximation by substituting $\theta_{\gamma}^{*}$ with an inexact estimate $\theta^{k}$.
@@
$\mathcal{G}(\theta^{k+1/2},x^{k},y^{k+1})\leq s_{k},\quad\text{or}\quad\mathcal{G}(\theta^{k+1/2},x^{k},y^{k+1})\leq\tau_{k}\mathcal{G}(\theta^{k-1},x^{k-1},y^{k-1}).$ (17)
@@
where $\nabla_{x}f\left(x^{k},\theta^{k+1/2}\right)+\nabla_{x}g\left(x^{k},\theta^{k+1/2}\right)$ is an approximation to $\nabla_{x}v_{\gamma}(x^{k},y^{k})=\nabla_{x}f(x^{k},\theta^{*}_{\gamma}(x^{k},y^{k}))+\nabla_{x}g(x^{k},\theta^{*}_{\gamma}(x^{k},y^{k}))$ (see (9)). Using the direction $d_{x}^{k}$, we update $x$ with stepsize $\alpha_{k}>0$ as
@@
1:  Input: Initial iterates $x^{0},y^{0},\tilde{\theta}^{0},\tilde{y}^{0}$, set $(x^{-1},y^{-1},\tilde{\theta}^{-1})=(x^{0},y^{0},\tilde{\theta}^{0})$, stepsizes $\alpha_{k},\beta_{k}$, relaxation parameter $\epsilon$, proximal parameter $\gamma$, penalty parameters $p_{0}$, $\varrho_{p}$, $c_{p}$, inexact parameters $\eta$, $s_{k}$, $\tau_{k}$, tolerance $tol$.
[…]
@@
9:        Case 1: $\Delta_{k+1}\geq c_{p}\min\big{\{}1/p_{k},\,t^{k+1}\big{\}}$;      Set $p_{k+1}:=p_{k}$, $\tilde{y}^{k+1}:=y^{k+1}$, $\tilde{\theta}^{k+1}:=\theta^{k+1}$.
10:        Case 2: $\Delta_{k+1}<c_{p}\min\big{\{}1/p_{k},\,t^{k+1}\big{\}},\,\|y^{k+1}-\theta^{k+1}\|\leq c_{y}\gamma/p_{k};$      Set $p_{k+1}:=p_{k}+\varrho_{p}$, $\tilde{y}^{k+1}:=y^{k+1}$, $\tilde{\theta}^{k+1}:=\theta^{k+1}$.
11:        Case 3: $\Delta_{k+1}<c_{p}\min\big{\{}1/p_{k},\,t^{k+1}\big{\}},\,\|y^{k+1}-\theta^{k+1}\|>c_{y}\gamma/p_{k};$      Apply the feasibility correction procedure (Algorithm 2).
@@ 4.1 Preliminary results for the analysis
Lemma 2.
Suppose Assumptions 2 and 3 hold, and let $\gamma>0$, $\eta>0$. Then, there exists $C_{\eta}>0$ such that for any $(x,y)\in X\times{\mathbb{R}}^{m}$, and $\theta\in{\mathbb{R}}^{m}$, it holds that
@@
Lemma 3.
Suppose Assumptions 2 and 3 hold. Let $\gamma\in(0,1/(\rho_{f_{2}}+\rho_{g_{2}}))$ and $(\bar{x},\bar{y})\in X\times{\mathbb{R}}^{m}$. Then, for any $(x,y)\in X\times{\mathbb{R}}^{m}$, the following inequalities hold,
@@
According to Proposition 2, for any $\bar{y}\in{\mathbb{R}}^{m}$, $v_{\gamma}(x,\bar{y})$ is a $(\rho_{f_{1}}+\rho_{g_{1}})$-weakly convex function with respect to $x$ on $X$, that is, $v_{\gamma}(x,\bar{y})+\frac{\rho_{f_{1}}+\rho_{g_{1}}}{2}\|x\|^{2}$ is convex on $X$. Then we can obtain (27) immediately. The second inequality (28) follows from Proposition 2 that for any $\bar{x}\in X$, $v_{\gamma}(\bar{x},y)$ is convex with respect to $y$. ∎
@@
Lemma 4.
Suppose Assumptions 2 and 3 hold. Then, there exists $C_{\gamma}>0$, such that the sequence $\{(x^{k},y^{k},\tilde{y}^{k})\}$ generated by AGILS (Algorithm 1) satisfies
@@
$\leq-\varphi(x^{k+1},\tilde{\theta}^{k+1})-\frac{1}{2\gamma}\|\tilde{\theta}^{k+1}-\tilde{y}^{k+1}\|^{2}+C_{\gamma}\zeta_{k+1}^{2}$
@@
Lemma 5.
Suppose Assumptions 1, 2 and 3 hold. Let $\gamma\in(0,1/(\rho_{f_{2}}+\rho_{g_{2}}))$, $\alpha_{k}\in[0,2/(L_{\psi_{x,k}}+c_{\alpha})]$ and $\beta_{k}\in[0,2/(L_{\psi_{y,k}}+c_{\beta})]$, where $L_{\psi_{x,k}}=L_{F_{x}}/p_{k}+L_{f_{x}}+L_{g_{1}}+\rho_{f_{1}}+\rho_{g_{1}}$, $L_{\psi_{y,k}}=L_{F_{y}}/p_{k}+L_{f_{y}}$ and $c_{\alpha},c_{\beta}$ are positive constants. Then, there exists $C_{\eta},C_{\gamma}>0$ such that the sequence $\{(x^{k},y^{k},\tilde{y}^{k})\}$ generated by AGILS (Algorithm 1) satisfies the following inequality,
@@
$=\frac{1}{\gamma}\left\|\theta^{*}_{\gamma}(x^{k},\tilde{y}^{k})-\theta^{k}\right\|$
$\leq\frac{C_{\eta}}{\gamma}\mathcal{G}({\theta}^{k},x^{k},\tilde{y}^{k})\leq\frac{C_{\eta}}{\gamma}\zeta_{k},$
where the first inequality follows from Lemma 2 and the second inequality is a consequence of the inexactness condition (20).
@@
where the last inequality follows from Lemma 2 and the inexact criterion (17). Combining the above inequality with (38) gives us
@@
Combining (35), (40), and then (29) established in Lemma 4, we obtain
@@
$-\left(\frac{1}{\alpha_{k}}-\frac{L_{\psi_{x,k}}}{2}-\frac{c_{\alpha}}{4}\right)\left\|x^{k+1}-x^{k}\right\|^{2}-\left(\frac{1}{\beta_{k}}-\frac{L_{\psi_{y,k}}}{2}-\frac{c_{\beta}}{4}\right)\left\|y^{k+1}-\tilde{y}^{k}\right\|^{2}.$

### v3

@@ 1 Introduction
where $\gamma>0$. When the lower-level problem is convex with respect to $y$, this reformulation is equivalent to the original bilevel optimization problem (1) (see [31, Theorem 1]). Directional optimality conditions for $({\rm VP})_{\gamma}$ are discussed in [5]. However, as shown in [31] and [62, Proposition 7], classical constraint qualification (CQ) conditions do not hold at any feasible point of $({\rm VP})_{\gamma}$, motivating us to study the following relaxed approximation problem,
@@ 2.1 Notations
Let $\mathbb{R}^{n}$ denote the $n$-dimensional Euclidean space, and let $\mathbb{B}$ denote the closed unit ball centered at the origin. The nonnegative and positive orthants in $\mathbb{R}^{n}$ are denoted by $\mathbb{R}^{n}_{+}$ and $\mathbb{R}^{n}_{++}$, respectively. The standard inner product and Euclidean norm are denoted by $\langle\cdot,\cdot\rangle$ and $\|\cdot\|$, respectively. For a vector $x\in\mathbb{R}^{n}$ and a closed convex set $S\subset\mathbb{R}^{n}$, the distance from $x$ to $S$ is defined as $\mathrm{dist}(x,S)=\min_{y\in S}\|x-y\|$. The indicator function $\delta_{S}:\mathbb{R}^{n}\to\mathbb{R}\cup\{+\infty\}$ is defined by $\delta_{S}(x)=0$ if $x\in S$ and $\delta_{S}(x)=+\infty$ otherwise. Moreover, $\mathcal{N}_{S}(x)$ denotes the normal cone to $S$ at $x$. The Euclidean projection operator onto $S$ is denoted by $\mathrm{Proj}_{S}$. The Cartesian product of two sets $X$ and $Y$ is denoted by $X\times Y$. Let $h:\mathbb{R}^{n}\to\mathbb{R}\cup\{\infty\}$ be a proper closed function. The proximal mapping of $h$ is defined as $\operatorname{Prox}_{h}(y):=\arg\min_{\theta\in{\mathbb{R}}^{m}}\left\{h(\theta)+\|\theta-y\|^{2}/2\right\}$. The Fréchet (regular) subdifferential $\hat{\partial}h$ and the Mordukhovich (limiting) subdifferential $\partial h$ at a point $\bar{z}\in\mathrm{dom}\,h:=\{z\mid h(z)<\infty\}$ are defined as:
$\hat{\partial}h(\bar{z})$ $=\left\{v\in\mathbb{R}^{n}\;\middle|\;h(z)\geq h(\bar{z})+\langle v,z-\bar{z}\rangle+o(\|z-\bar{z}\|),\;\forall z\in\mathbb{R}^{n}\right\},$
@@ 2.2 Basic Assumptions
Remark 2.4.
The bilevel problem (1) has also been studied in [42]. However, in addition to Assumption 2.3, Assumption 3.2 in [42] imposes an additional Lipschitz-like continuity condition on the proximal operator $\operatorname{Prox}_{{g}(x,\cdot)}$ with respect to $x$. This assumption is challenging to verify in practice and is not required in this work.
Assumption 2.3(1) implies $(\rho_{f_{1}},\rho_{f_{2}})$-weak convexity with $\rho_{f_{1}}=\rho_{f_{2}}=L_{f}$ of $f(x,y)$ [7, Lemma 5.7]. Furthermore, the lower-level objective $\varphi(x,y)$ is $(\rho_{\varphi_{1}},\rho_{\varphi_{2}})$-weakly convex on $X\times Y$, where $\rho_{\varphi_{1}}=\rho_{f_{1}}+\rho_{g_{1}}$ and $\rho_{\varphi_{2}}=\rho_{f_{2}}+\rho_{g_{2}}$.
According to [24, Lemma 4.2], $g(x,y)=h(c(x,y))$ is $L_{h}L_{c}$-weakly convex if $h:\mathbb{R}^{d}\to\mathbb{R}$ is a closed convex $L_{h}$-Lipschitz continuous function and $c:\mathbb{R}^{n+m}\to\mathbb{R}^{d}$ is a smooth function with an $L_{c}$-Lipschitz continuous Jacobian. Below, we provide more specific examples of $g(x,y)$ satisfying Assumption 2.3.
Example 2.5.
Consider $g(x,y)=xP(y)$, where $P(y)$ is a convex and $L_{P}$-Lipschitz continuous function on $Y$. Then $g(x,y)$ is $(\rho_{g_{1}},\rho_{g_{2}})$-weakly convex on $\mathbb{R}_{+}\times\mathbb{R}^{m}$ for any $\rho_{g_{1}},\rho_{g_{2}}\geq 0$ satisfying $\rho_{g_{1}}\rho_{g_{2}}\geq L_{P}^{2}$.
@@
(1) Let $g(x,y)=x\|y\|_{1}$. Then $g(x,y)=xP(y)$ with $P(y)=\|y\|_{1}$, which is $\sqrt{m}$-Lipschitz on $\mathbb{R}^{m}$. Thus, $g(x,y)$ is $(m,1)$-weakly convex on $\mathbb{R}_{+}\times\mathbb{R}^{m}$.
(2) Let $g(x,y)=\sum_{j=1}^{J}x_{j}\|y^{(j)}\|_{2}$, where $y^{(j)}$ denotes the $j$-th group of $y$. Then $g(x,y)=\sum_{j=1}^{J}x_{j}P_{j}(y^{(j)})$ with $P_{j}(y^{(j)})=\|y^{(j)}\|_{2}$, which is $1$-Lipschitz. Thus, $g(x,y)$ is $(1,1)$-weakly convex on $\mathbb{R}^{J}_{+}\times\mathbb{R}^{m}$.
@@ 2.3 On the Moreau Envelope Function Reformulation
Under Assumption 2.2, the solution mapping
@@
Proposition 2.7.
Suppose Assumptions 2.2 and 2.3 hold. For $\gamma\in(0,1/(2\rho_{f_{2}}+2\rho_{g_{2}}))$, the function $v_{\gamma}(x,y)$ is $(\rho_{v_{1}},\rho_{v_{2}})$-weakly convex on $X\times\mathbb{R}^{m}$ with $\rho_{v_{1}}\geq\rho_{f_{1}}+\rho_{g_{1}}$ and $\rho_{v_{2}}\geq 1/\gamma$. Additionally, $v_{\gamma}(x,y)$ is differentiable on $X\times\mathbb{R}^{m}$ with gradient
@@
Proposition 2.8.
Suppose Assumptions 2.2 and 2.3 hold. For $\gamma\in(0,1/(\rho_{f_{2}}+\rho_{g_{2}}))$, the function $v_{\gamma}(x,y)$ is $\rho_{v_{1}}$-weakly convex with respect to $x$ on $X$ for any $y\in Y$ with $\rho_{v_{1}}\geq\rho_{f_{1}}+\rho_{g_{1}}$. It is also convex with respect to $y$ on $\mathbb{R}^{m}$ for any $x\in X$. Moreover, $v_{\gamma}(x,y)$ is differentiable with gradient given by (9) on $X\times\mathbb{R}^{m}$, and $\theta_{\gamma}^{*}(x,y)$ and $\nabla v_{\gamma}(x,y)$ are continuous on $X\times\mathbb{R}^{m}$.
@@
We begin by showing that $v_{\gamma}(x,y)$ is continuously differentiable on $X\times\mathbb{R}^{m}$. By Assumption 2.2, the function $f(x,\theta)+g(x,\theta)+\frac{1}{2\gamma}\|\theta-y\|^{2}$ is level-bounded in $\theta$ locally uniformly for any $(\bar{x},\bar{y})\in X\times\mathbb{R}^{m}$. Specifically, for any $c\in\mathbb{R}$, there exist a compact set $D$ and a neighborhood $Z$ of $(\bar{x},\bar{y})$ such that the level set $\{\theta\in\mathbb{R}^{m}\mid f(x,\theta)+g(x,\theta)+\frac{1}{2\gamma}\|\theta-y\|^{2}\leq c\}$ is contained in $D$ for all $(x,y)\in Z$. Using Corollary 4.3(ii) from [49], it follows that $v_{\gamma}(x,y)$ is locally Lipschitz continuous on $X\times\mathbb{R}^{m}$. Furthermore, Theorem 1.22, Theorem 4.1(ii) and Theorem 4.17 in [49] imply that $v_{\gamma}({x},{y})$ is differentiable on $X\times\mathbb{R}^{m}$ with its gradient given as in (9). The partial weak convexity property of $v_{\gamma}({x},{y})$ can be established using arguments similar to those in the proof of Theorem 2 in [31].
@@
Lemma 2.9.
Suppose Assumptions 2.2 and 2.3 hold. The mapping $\theta_{\gamma}^{*}(x,y)$ is $L_{\theta^{*}}$-Lipschitz continuous with respect to $y$ for any $x\in X$, with $L_{\theta^{*}}:=1+\sqrt{1-s/\gamma}$ where $s=\gamma/(\gamma L_{f_{y}}+1)^{2}$.
@@
$0\in\nabla_{y}f(x,\theta^{*}_{i})+\partial_{y}g(x,\theta^{*}_{i})+(\theta^{*}_{i}-y_{i})/\gamma+\mathcal{N}_{Y}(\theta^{*}_{i}),\quad\mathrm{for}~i=1,2.$
@@ 3 Algorithm design
where $p_{k}$ is the penalty parameter at iteration $k$. Due to the nonsmoothness of $g(x,y)$ with respect to $y$ and the availability of its proximal operator, it is natural to adopt a proximal alternating linearization method [14] to construct the algorithm. Given the current iterate $(x^{k},y^{k})$, linearizing the smooth components of $\psi_{p_{k}}(x,y)$ requires evaluating gradients, including $\nabla v_{\gamma}(x,y)$ as provided in Proposition 2.7. Calculating the exact gradient $\nabla v_{\gamma}(x^{k},y^{k})$ requires solving the proximal lower-level problem in (8) to obtain $\theta_{\gamma}^{*}(x^{k},y^{k})$. However, this solution typically does not have a closed-form expression, and solving it requires an iterative optimization solver, which can be computationally expensive. To address this, we propose an inexact gradient approximation by substituting $\theta_{\gamma}^{*}$ with an inexact estimate $\theta^{k}$.
@@
$\mathcal{G}(\theta^{k+1/2},x^{k},y^{k+1})\leq s_{k},\quad\text{or}\quad\mathcal{G}(\theta^{k+1/2},x^{k},y^{k+1})\leq\tau_{k}\mathcal{G}(\theta^{k-1},x^{k-1},y^{k-1}),$ (17)
which can be compactly expressed as
$\mathcal{G}(\theta^{k+1/2},x^{k},y^{k+1})\leq\max\left\{s_{k},\ \tau_{k}\,\mathcal{G}(\theta^{k-1},x^{k-1},y^{k-1})\right\}.$
@@
where $\nabla_{x}f\left(x^{k},\theta^{k+1/2}\right)+\nabla_{x}g\left(x^{k},\theta^{k+1/2}\right)$ is an approximation to $\nabla_{x}v_{\gamma}(x^{k},y^{k+1})=\nabla_{x}f(x^{k},\theta^{*}_{\gamma}(x^{k},y^{k+1}))+\nabla_{x}g(x^{k},\theta^{*}_{\gamma}(x^{k},y^{k+1}))$ (see (9)). Using the direction $d_{x}^{k}$, we update $x$ with stepsize $\alpha_{k}>0$ as
@@
1:Input: Initial iterates $x^{0},\theta^{-1},y^{0},\tilde{\theta}^{0},\tilde{y}^{0}$, set $(x^{-1},y^{-1},\theta^{-1})=(x^{0},y^{0},\tilde{\theta}^{0})$, stepsizes $\alpha_{k},\beta_{k}$, relaxation parameter $\epsilon$, proximal parameter $\gamma$, penalty parameters $p_{0}$, $\varrho_{p}$, $c_{p}$, $c_{y}$, inexact parameters $\eta$, $s_{k}$, $\tau_{k}$, tolerance toltol.
2:for $k=0,\ 1,\cdots$ do
@@
9:   Case 1: $\Delta_{k+1}\geq c_{p}\min\big\{1/p_{k},\,t^{k+1}\big\}$      Set $p_{k+1}:=p_{k}$, $\tilde{y}^{k+1}:=y^{k+1}$, $\tilde{\theta}^{k+1}:=\theta^{k+1}$.
10:   Case 2: $\Delta_{k+1}<c_{p}\min\big\{1/p_{k},\,t^{k+1}\big\},\,\|y^{k+1}-\theta^{k+1}\|\leq c_{y}\gamma/p_{k}$     Set $p_{k+1}:=p_{k}+\varrho_{p}$, $\tilde{y}^{k+1}:=y^{k+1}$, $\tilde{\theta}^{k+1}:=\theta^{k+1}$.
11:   Case 3: $\Delta_{k+1}<c_{p}\min\big\{1/p_{k},\,t^{k+1}\big\},\,\|y^{k+1}-\theta^{k+1}\|>c_{y}\gamma/p_{k}$     Apply the feasibility correction procedure (Algorithm 2).
@@ 4.1 Preliminary results for the analysis
Lemma 4.1.
Suppose Assumptions 2.2 and 2.3 hold, and let $\gamma>0$, $\eta>0$. Then, there exists $C_{\eta}>0$ such that for any $(x,y)\in X\times{\mathbb{R}}^{m}$, and $\theta\in{\mathbb{R}}^{m}$, it holds that
@@
Lemma 4.2.
Suppose Assumptions 2.2 and 2.3 hold. Let $\gamma\in(0,1/(\rho_{f_{2}}+\rho_{g_{2}}))$ and $(\bar{x},\bar{y})\in X\times{\mathbb{R}}^{m}$. Then, for any $(x,y)\in X\times{\mathbb{R}}^{m}$, the following inequalities hold,
@@
According to Proposition 2.8, for any $\bar{y}\in{\mathbb{R}}^{m}$, $v_{\gamma}(x,\bar{y})$ is a $(\rho_{f_{1}}+\rho_{g_{1}})$-weakly convex function with respect to $x$ on $X$, that is, $v_{\gamma}(x,\bar{y})+\frac{\rho_{f_{1}}+\rho_{g_{1}}}{2}\|x\|^{2}$ is convex on $X$. Then we can obtain (27) immediately. The second inequality (28) follows from Proposition 2.8 that for any $\bar{x}\in X$, $v_{\gamma}(\bar{x},y)$ is convex with respect to $y$. ∎
@@
Lemma 4.3.
Suppose Assumptions 2.2 and 2.3 hold. Then, there exists $C_{\gamma}>0$, such that the sequence $\{(x^{k},y^{k},\tilde{y}^{k})\}$ generated by AGILS (Algorithm 1) satisfies
@@
$\leq-\varphi(x^{k+1},\tilde{\theta}^{k+1})-\frac{1}{2\gamma}\|\tilde{\theta}^{k+1}-\tilde{y}^{k+1}\|^{2}+C_{\gamma}\zeta_{k+1}^{2}.$
@@
Lemma 4.4.
Suppose Assumptions 2.1, 2.2 and 2.3 hold. Let $\gamma\in(0,1/(\rho_{f_{2}}+\rho_{g_{2}}))$, $\alpha_{k}\in[0,2/(L_{\psi_{x,k}}+c_{\alpha})]$ and $\beta_{k}\in[0,2/(L_{\psi_{y,k}}+c_{\beta})]$, where $L_{\psi_{x,k}}=L_{F_{x}}/p_{k}+L_{f_{x}}+L_{g_{1}}+\rho_{f_{1}}+\rho_{g_{1}}$, $L_{\psi_{y,k}}=L_{F_{y}}/p_{k}+L_{f_{y}}$ and $c_{\alpha},c_{\beta}$ are positive constants. Then, there exists $C_{\eta},C_{\gamma}>0$ such that the sequence $\{(x^{k},y^{k},\tilde{y}^{k})\}$ generated by AGILS (Algorithm 1) satisfies the following inequality,
@@
$=\frac{1}{\gamma}\left\|\theta^{*}_{\gamma}(x^{k},\tilde{y}^{k})-\theta^{k}\right\|\leq\frac{C_{\eta}}{\gamma}\mathcal{G}({\theta}^{k},x^{k},\tilde{y}^{k})\leq\frac{C_{\eta}}{\gamma}\zeta_{k},$
where the first inequality follows from Lemma 4.1 and the second inequality is a consequence of the inexactness condition (20).
@@
where the last inequality follows from Lemma 4.1 and the inexact criterion (17). Combining the above inequality with (38) gives us
@@
Combining (35), (40), and then (29) established in Lemma 4.3, we obtain
@@
$\hskip-50.0pt-\left(\frac{1}{\alpha_{k}}-\frac{L_{\psi_{x,k}}}{2}-\frac{c_{\alpha}}{4}\right)\left\|x^{k+1}-x^{k}\right\|^{2}-\left(\frac{1}{\beta_{k}}-\frac{L_{\psi_{y,k}}}{2}-\frac{c_{\beta}}{4}\right)\left\|y^{k+1}-\tilde{y}^{k}\right\|^{2}.$